Shift summary for evening shift, 8/18/99

Christine Pietrykowski (pietryko@hep.uchicago.edu)
Thu, 19 Aug 1999 00:26:08 -0500

Shift Summary for evening shift, 08/18/99

On Shift: CP,LB

Hours of beam available at all intensities:
Hours of beam available at nominal intensity:3:59 hr.
Hours of E832 data (special runs not included):3:53 hr.
Hours of special runs (laser scans, muon runs...):8 min.

Accounting of lost time:
Linac Klystron/RF problems
Problems with NM2EU PS
========================================================================
Runs ended this shift:
14259 BEAM_E832
14260 BEAM_MUON
14261 BEAM_E832 (NM2EU problems)
14263 BEAM_E832 (short)
========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
E832 data taken when possible.
========================================================================
Description of Shift:
run spill time comment
------------------------------------------------------------------------
14259 56 16:42 Beam went away. LINAC klystron problems, experts
enroute. No estimate for beam.
125 18:08 Linac RF problems continue. Still no beam. End
Run.
14260 18:18 Set-up for a Muon Run with a polarity change.
43 19:16 Beam back up. Start Muon Run.
47 19:24 Terminate Muon run.
14261 19:36 Start Hadron run.
21:20 Problems with the NM2EU PS, beam is
not well located on target. Questionable data.
21:47 Called MCR to ask that the beamstop be closed,
out of concern for DC currents. Technitions on
scene working on NM2EU PS.
98 21:48 Terminate Run.
14262 22:30 Run COSMICS
22:50 End COSMICS
14263 22:59 Start Hadron run.
4 23:04 Improved beam quality being delivered.
24 23:31 End run, DDD was hung at plane 1, stream 3.
Reset. Working on a dart-fresh-start.
23:37 Lost beam. NM2EU PS problems again. Electricians
on the scene.
23:40 Problems with Air Conditioning unit in counting
room. AC-3; problem is old loose belts, to be
fixed in the morning.
23:45 Phone call from Rick Ford, said that failure of
NM2EU began about 24 hours ago, although it was
intermittent. Deterioration was worse starting
at 5:30 - 6:00PM, shortly before it was detected.
Rick Kessler investigating possiblity of
identifying good spills in this time frame using
SDAQ SWIC database.
========================================================================
Items to Follow up on Next Shift:
NM2EU PS problem
Air Conditioning unit
~